# Towards Autonomous Driving with Small-Scale Cars: A Survey of Recent Development

### v1

@@ II-C Sim2Real transfer
The first approach for the real-world application of small-scale cars is Zero-shot Transfer, where the trained model is directly applied in real-world settings. For this, the Imitation Learning (IL) approach is usually involved with the dataset collected in the real-world platform. The agent is trained to mimic the behavior of the expert [34, 90, 9, 56, 91, 55, 92, 65]. In addition to IL, Zero-shot Transfer for DRL can be achieved with compact observation and output space [93, 94, 95]. Another noteworthy approach is Transfer Learning, which aims to improve the performance of the target agent in the target domain by transferring the knowledge contained in different but related source domains [96], for example, for Sim2Real transfer, the process typically includes initially pre-training a model in a simulated environment and subsequently fine-tuning it using real-world data. When the labeled data in the target domain is scarce or expensive to obtain, Domain Adaptation is used. As a subset of transfer learning, it seeks specifically to minimize the distribution mismatch between the source and target domains, allowing the model to generalize better across different domains. Another frequently used approach is Domain Randomization, in which the randomness and variation are introduced during simulation to make the model more robust to different real-world conditions [97]. It works by for example randomly varying simulation parameters, such as lighting conditions, textures, object appearances, and physics properties during training. For small-scale car platforms, Duckietown and DeepRacer provide Domain Randomization option in the simulation, which is easier for the users to tackle Sim2Real issue [98, 10, 99, 38].
@@ III Benchmark Tasks
Autonomous driving systems for normal and small-scale vehicles are normally categorized into two distinct pipelines: the end-to-end system pipeline and the modular system pipeline. As illustrated in Fig. 2, a modular system comprises multiple subsystems that perform various tasks such as perception and localization, mapping, path planning, and control [100, 101]. Each subsystem focuses on specific functionalities and tasks. First is the perception system, where sensors such as cameras and LiDAR are used to gather information about the surroundings, including lane markings, obstacles, traffic signs, and other vehicles. Following this, the localization and mapping system leverages GPS, odometry, IMU, or techniques like simultaneous localization and mapping (SLAM) to precisely pinpoint the position within the environment while concurrently creating a detailed map of its surroundings. The perception system creates an intermediate representation of the environment for subsequent utilization.
@@
The representation module then uses the information from the perception module and further processes the sensor data with sensor fusion techniques or creates an object map with the predicted state of each object within the sensor range. The combination of the perception module and the representation module can be seen as the scene understanding, which provides an abstract high-level representation of the environment [101]. Afterward, the planning system maps out a safe and efficient route to reach the destination. Normally, in the autonomous driving system, the planning phase is divided into two different parts, namely global path planning and local path planning [102]. Global path planning refers to the process of determining an optimal or feasible route from the current position to its destination and is done considering the entire environment and involves high-level decision-making. Local path planning or path following, on the other hand, focuses on the immediate surroundings of the vehicle and deals with making real-time adjustments to adhere to the planned global path. Finally, the control module generates driving commands based on the processed information. In control theory, the primary goal is to minimize a cost function. Various methodologies are employed for the control system, classical controllers such as the Proportional–Integral–Derivative (PID) controller, Model Predictive Control (MPC), and ML-based controllers such as IL or DRL.
@@
An end-to-end system, on the other hand, is characterized by a unified architecture that aims to learn the entire mapping directly from raw sensor inputs to driving actions without explicitly decomposing the task into separate modules [103, 104, 105]. It involves training a comprehensive learning model with ML methods, IL or DRL, directly process raw sensor data and output control commands. End-to-end systems potentially simplify the system architecture by eliminating the need for handcrafted modules and feature engineering.
In this section, our focus is directed toward discussing various tasks studied in the literature and the methods developed to handle these tasks using modular or end-to-end systems. The detailed methodologies for achieving these tasks will not be discussed, instead, we will refer to the respective papers. The benchmark tasks categorized in this survey are illustrated in Fig. 3.
Figure 3: Research in the literature concerning small-scale cars is classified into various benchmark tasks in this survey, ranging from localization and mapping, path planning and following, lane keeping to racing, and cooperative driving.
III-A Localization and Mapping
The localization and mapping, which, although commonly viewed as a subsystem within the modular autonomous system, are also regarded as distinct and extensively researched topics in the literature.

### v2

@@ II-C Sim2Real Transfer
The first approach for the real-world application of small-scale cars is Zero-shot Transfer, where the trained model is directly applied in real-world settings. Imitation Learning (IL), based on datasets from real platforms, trains agents to mimic expert behavior [33, 81, 11, 55, 82, 54, 83, 62]. While efficient, this method assumes simulation and real-world environments are similar. DRL-based Zero-Shot Transfer, achieved via compact observation and output spaces [84, 85, 86], depends heavily on simulation fidelity for success. Another noteworthy approach is Transfer Learning, which aims to improve the performance of the target agent in the target domain by transferring the knowledge contained in different but related source domains [87]. It excels when simulation closely aligns with real-world scenarios but falters with significant domain mismatches. For small-scale cars, this method balances efficiency and adaptability in controlled conditions. When the labeled data in the target domain is scarce or expensive, Domain Adaptation is used. As a subset of transfer learning, it seeks specifically to minimize the distribution mismatch between the source and target domains, allowing the model to generalize better across different domains. It is effective in dynamic environments but requires computational resources and high-quality feature extraction, posing challenges in real-time applications. Another frequently used approach is Domain Randomization, in which the randomness and variation are introduced during simulation to make the model more robust to different real-world conditions [88]. It works by for example randomly varying simulation parameters, such as lighting conditions, textures, object appearances, and physics properties during training. For small-scale car platforms, Duckietown and DeepRacer provide Domain Randomization option in the simulation, which is easier for the users to tackle Sim2Real issue [89, 10, 90, 37]. While fostering generalization, excessive randomization may dilute the relevance of training data, compromising real-world performance.
III Sensors and Sensor systems
In Section II-A, various platforms are explored, each employing diverse sensors to perceive their surroundings and execute tasks. Compared with full-scale AVs, which are equipped with various state-of-the-art sensors for perception [91], small-scale cars usually use lightweight sensors. Here, we discuss the most commonly utilized sensors across these platforms, providing their respective use cases.
III-A Camera
An RGB camera provides a stream of images that is easily processed and understood by humans. This helps researchers to better understand the perspective of the robot. The amount of information provided by cameras is huge, which allows the robot to perceive a lot of information with just one sensor in a short amount of time. The downside of this is, that they need a high bandwidth communication link to whichever part of the robot needs those information. Also with the information in images being so densely encoded, it is difficult for a robot to understand the environment.
The field of computer vision has emerged as a crucial scientific discipline, facilitating the extraction of pertinent details from images and videos. In recent years, the advent of deep learning (DL) models, such as You Only Look Once (YOLO) [92], has significantly contributed to enhancing the robot’s ability to interpret scenes captured by cameras. It is worth noting that cameras face external influences, such as varying lighting conditions, which can substantially alter the captured images. In extreme scenarios, such as very dark environments or instances where direct light affects the camera, the effectiveness of cameras may be compromised. A big advantage of cameras is their wide pricing range, starting from 15 USD, corresponding to images of different quality in regards to resolution and dynamic range. They are also easy to interface with, usually requiring only a USB connection. Some tasks can be accomplished only with cameras, e.g. traffic sign or road marking detection [93, 94, 95].
Beyond RGB cameras, more advanced options exist, including those detecting optical flow [96] (the apparent motion of objects) or depth cameras [97], employing different methods to ascertain the distance to objects. These advanced cameras provide both depth information and standard RGB images. In the context of autonomous tasks for single-car platforms discussed in this paper, cameras emerge as a ubiquitous choice. Their cost-effectiveness and information-rich output make them the preferred sensor for many platforms. Among the commonly used cameras in scaled vehicles, the Raspberry Pi Camera stands out [9, 23, 31, 62, 28, 59]. Designed for the Raspberry Pi, it combines affordability with decent image quality, offering a resolution of 1080p and a dynamic range of 44dB. Another popular option used in the discussed platforms is the Intel Realsense line of cameras [25, 26]. Those include two cameras and an IR projector, which allow to capture not only an RGB image but also depth information of each pixel using stereovision. The IR projector helps to improve accuracy in scenes with poor textures. The manufacturer claims an accuracy in the centimeter range.
The big advantage of cameras is their versatility combined with their low price. Their dense stream of information makes them suitable for almost all tasks discussed in this paper. Although different approaches for processing the information from cameras need to be applied depending on the task. Applying DL approaches helps with this in many different fields and allows for quickly iterating through different approaches. One major drawback of relying solely on cameras is their susceptibility to varying environmental conditions, as previously mentioned. This sensitivity complicates the transfer of control strategies from simulations to real-world vehicles. For successful transfer, simulations must either model these environmental influences with high accuracy or incorporate an additional interpretation layer that bridges the gap between simulation and reality.
III-B LiDAR
LiDAR, another widely employed sensor in numerous platforms, functions by measuring the time taken for a beam of light to reflect off a surface. The LiDAR emits beams from multiple angles—typically ranging from 360 to 2048—creating a detailed point cloud map of objects in the environment. Multi-layer LiDAR systems, with laser beams spanning up to 128 angles, provide data on objects at different heights, generating a comprehensive 3D point cloud. LiDARs are available in various resolutions, offering enhanced detail, especially in complex and structured environments. The typical scanning frequency of LiDARs ranges from 5 to 20 Hz, with high-end models achieving detection ranges of up to 350 meters.
The distinct advantage of LiDAR over cameras lies in its direct identification of the surroundings, including the precise position of objects relative to the robot. In contrast, a camera first needs to identify an object in an image and then estimate its position. Moreover, LiDAR provides significantly more accurate positional information compared to estimations derived from camera images. However, unlike cameras, LiDARs lack the ability to discern details such as the color of the detected objects. This limitation renders them unsuitable, for example, in tasks like traffic signs or lane detection. LiDAR also has a significant advantage over cameras in that it is unaffected by environmental conditions such as varying lighting, and it can operate effectively in complete darkness. Additionally, some LiDAR models offer reflectivity data, which can help distinguish between different surface types. For example, they can differentiate highly reflective road markings from less reflective obstacles. However, LiDARs are typically more expensive than cameras. The most affordable hobbyist LiDARs start at around 150 USD for a basic single-plane sensor, with prices increasing rapidly as resolution or the number of planes expands.
Usually physically larger platforms use LiDAR sensors [10, 23, 12, 24, 13, 25, 26, 62, 27, 59, 18] as they tend to be bigger and heavier than cameras. Those platforms are also usually more versatile and used for more different tasks and therefore the additional cost of a LiDAR is feasible. In our investigation, all platforms utilizing LiDAR also integrate a camera to address the limitations of LiDAR for specific tasks. The LIDAR that is used by far the most on the platforms we investigated is the single plane RPLIDAR [23, 13, 26, 62, 27, 59]. It offers a reasonable detection range of 12m and a resolution of 0.225 degrees for small-scale car platforms with a 360 field of view. And is one of the cheaper options costing around 400 USD. But it only allows for a 10 Hz scan frequency. The second most popular, higher end solution is the Hokuyo UST-10LX [12, 24], offering a 270 degree field of view and a 10m detection range.
Particularly tasks, that require precise information about the location and orientation of surrounding vehicles or obstacles benefit greatly from the improved accuracy of LIDARs over cameras, these include car following, overtaking, or obstacle avoidance. Tasks such as lane following or traffic sign following need the additional visual information provided by cameras and can therefore not be accomplished with only LIDARs.
For economic considerations, cost-effective sensors are the mainstream choice for small-scale cars. These sensors typically offer lower resolution but demand less computational power. For example, the RPLiDAR is commonly used in small-scale cars, providing a detection range of up to 40 meters with error margins of approximately 1%. It has a data rate of around 32,000 samples per second. In contrast, typical LiDAR sensors used in full-scale AVs offer significantly higher range, resolution, and data rates. For instance, the Velodyne LiDAR sensor can detect up to 120 meters with a 2 cm error margin and a data rate of 1.3 million points per second. These hardware limitations make it challenging to develop new approaches, leading primarily to the application of mature AV techniques from full-scale cars under simplified conditions. However, advancements in semiconductor technology are expected to make more sophisticated hardware setups available for small-scale cars in the near future, which will unlock more advanced AV techniques for small-scale cars.
III-C IMU and Encoder
The information of these two types of sensors is usually combined with other absolute positioning inputs like a camera or a central localization system to fuse them in a filter like a Kalman filter [98] or in any other kind of perception module. IMUs are available at relatively affordable prices, ranging from 15 USD. The two IMU sensors most commonly used in the discussed platforms are the MPU9250 and MPU6050 [9, 23, 17, 26, 59]. Although it has to be noted, that not all publications mention the exact model of the IMU used. Both sensors feature a 3-axis gyroscope, which measures the rotational speed, and a 3-axis accelerometer, which measures linear acceleration. The MPU9250 additionally offers a 3-axis magnetometer to incorporate measurements of the Earth’s magnetic field into orientation sensing.
Odometry encoders can be found in several different forms, ranging from magnetic encoders to optical ones. Their implementation is usually dependent on the platform, as they need to be tightly integrated into the mechanics of the robot. Despite their low cost, often under 10 USD, they offer reliable positional data, making them a popular choice. Almost all the platforms reviewed in this paper incorporate an IMU and some form of encoder, as these sensors are both affordable and effective, particularly in slow, controlled indoor environments. The notable exception is the smallest platform, primarily designed for swarm robotics research [22], where extreme cost constraints led to the exclusion of these sensors.
The addition of cheap odometry encoders or IMUs especially benefits tasks that require the knowledge of the precise location and orientation of the vehicles. This includes tasks such as path following, racing, or drifting. The disadvantage of this type of sensor lies in the need to incorporate it in a meaningful way, this usually involves either existing filters or using deep neural networks. In the first case, the additional filter needs to be tuned to provide meaningful information, in the case of the latter the training of the neural network becomes more complex. The big advantage lies in the small price and easily accessible information.
III-D GPS
GPS is the least frequently used sensor with model scale platforms [11, 13] offer positional information in a global coordinate system. Most of the discussed platforms are designed for indoor use, which eliminates the use of GPS as the signal does not penetrate buildings. However, recognizing the pivotal role GPS plays in full-sized vehicles, many model-scale platforms opt to replace it with indoor central localization systems, as detailed in Section III-E, to enhance realism. GPS receivers are also comparatively cheap at around 30 USD. However, for those seeking higher precision, the cost can escalate to several hundred USD, especially for the more advanced GPS RTK systems [99]. Such precision may become necessary for model-scale platforms, as conventional GPS systems typically only offer an accuracy of approximately 5 meters [100]. The biggest disadvantage of GPS sensors for small-scale vehicles is their reliance on outdoor environments, which require significantly more space and make experiments heavily dependent on weather and other outdoor conditions. As a result, GPS sensors are only beneficial for platforms when the research focuses specifically on the accurate effects and phenomena associated with GPS data, such as multipath errors or signal interference.
III-E Central Localization System
In situations where the emphasis is on studying the interaction among multiple vehicles rather than individual vehicle behavior, a central localization system streamlines the task, allowing a more focused approach to other aspects such as obstacle avoidance or overtaking, by providing precise information about the position and orientation of different objects and sometimes even their velocity. By decoupling the algorithms for (multi) vehicle behavior from the perception of individual vehicles, central systems enable the study of system dynamics in an idealized ”best case” scenario. In this context, the term ”best case” refers to a situation where a vehicle possesses precise knowledge of its own position and that of surrounding obstacles. However, it is essential to acknowledge that this assumption may not align with the complexities of real-world road situations. Conversely, central systems find utility in simulating GPS positioning within actual vehicles.
One of the cheapest options implemented in the considered platforms involves the use of special markers or tags on each vehicle, which are observed by one or multiple overhead cameras [101]. These markers, based on ArUco [102] technology, are strategically placed on vehicles and key positions or obstacles. Knowing the intrinsic and extrinsic parameters of the cameras, existing libraries like OpenCV [103] can be employed to detect these markers and estimate their position and orientation. As discussed in Section III-A, cameras are relatively inexpensive, and the markers merely need to be printed out. While the resulting localization precision is reasonable [104], it does decline with the distance of the markers from the camera. Mitigating this issue involves the use of multiple cameras, necessitating synchronization and precise positioning of all cameras.
Off-the-shelf solutions for tracking also exist from various manufacturers, such as OptiTrack or VICON. These solutions predominantly rely on high-speed IR cameras and reflective markers on the tracked objects. The systems inherently provide position and orientation information for different marker assemblies (referred to as rigid bodies) without requiring additional processing. Notably, the accuracy of these systems surpasses that of solutions utilizing RGB cameras, with manufacturers often specifying sub-millimeter accuracy. However, the upfront cost is relatively high, typically exceeding 10,000 USD, and in some cases, additional licensing fees for the requisite software may apply [63, 21].
Central localization systems offer the advantage of high precision localization of any object within the test setup and are easily extendable to more vehicles of objects as no additional sensors need to be mounted on the vehicles themselves. They are however quite expensive and require a dedicated mounting setup to position the sensors, therefore they cannot easily be relocated. Their high precision allows for easy transferability from simulation to real-world testbed. The implementation of these systems especially benefits tasks, that do not study the influence of sensors and sensor behavior on the control algorithm, but merely the control strategies themselves. They can essentially be used as the single source of information for any of the tasks listed except traffic sign following.
III-F Other sensors
Several platforms discussed in Section II-A employ additional sensors to either emulate sensors found in full-scale autonomous cars or fulfill specific tasks.
A prevalent type of sensor utilized in many platforms [9, 17, 30, 32, 26, 62, 64, 66, 28, 105] is the single-point distance sensor, available as either ultrasonic or infrared. These sensors operate on the time-of-flight principle. They measure the time it takes for a wave to reflect back to the sensor and calculate the distance accordingly. Typically positioned at the front of vehicles, they often mimic radar sensors and play a crucial role in tasks such as car following.
Line following sensors, consisting of an array of light sensors aimed at the ground, are employed in certain platforms [26, 21, 26] to assist visual line following, with a camera serving as a secondary input. These sensors can easily discern the line from the road by measuring the reflectance of the surface on multiple points along the length of the sensor. A limitation is their coverage area, which is relatively small and located underneath the car. If the vehicle deviates significantly from the line, the sensor may struggle to rediscover it.
One platform [101] utilizes RFID readers to detect specific points of interest on the road, such as the beginning of intersections or as indicators for turns on a crossroad. Turtlebot employs cliff sensors, which function as downward-facing distance sensors, indicating if the part of the vehicle with the sensor is suspended over a steep drop. This sensor serves to prevent the vehicle from inadvertently falling off a precipice in the environment.
For the platform designed for use with a high amount of robots [22], infrared (IR) transmitters and receivers are employed. These facilitate cost-effective communication between different robots and emulate a simple form of Vehicle-to-Vehicle (V2V) communication.
III-G Compute Units
To process the information from the sensors and compute actuator movement according to their task, the platforms require a dedicated computational unit. As the physical size of the platforms is quite limited, the computational performance of these compute units is as well. The prevailing choice across numerous platforms [9, 23, 26, 21, 63, 106, 60, 101, 28] is the Raspberry Pi single-board computer (SBC). It offers a performance of about 32 GFLOPS leveraging both CPU and GPU. Built on a System on a Chip (SoC), the Raspberry Pi features a compact physical size but offers modest computational performance, with a 4-core 1.5GHz CPU. Despite its relatively low processing power, the Raspberry Pi is well-suited for the majority of tasks performed by these platforms. The Raspberry Pi also benefits from a large developer and research community providing a diverse set of frameworks and libraries to harness its capabilities.
A comparable alternative to the Raspberry Pi is the Nvidia Jetson. Also based on an SoC, the Nvidia Jetson stands out by incorporating an additional higher performance GPU to accelerate ML tasks, resulting in a performance of around 472 GFLOPS. Different versions of the Nvidia Jetson are available, some also offering dedicated resources for tensor operations for DL tasks. The highest performing option offers up to 300 TOPS. Given their similar connectivity, some platforms offer compatibility with either the Raspberry Pi or the Nvidia Jetson [9, 23]. However, the majority of platforms exclusively support the Nvidia Jetson [12, 24, 25, 18, 59, 65].
For platforms not requiring high computational performance on the vehicles themselves, microcontrollers [17, 30, 61, 22, 67, 64] are very common. Although microcontrollers offer lower performance than Raspberry Pi, offering only between 100 KFLOPS and 25 MFLOPS, they enable real-time code execution, crucial for interfacing with low-level sensors and actuators. They also use less power and are cheaper. Conversely, some platforms leverage regular PCs as their computational unit [10, 11]. While PCs provide the highest performance, ranging from 50 GFLOPS for CPU-only tasks up to multiple TFLOPS for GPU computation, they have larger physical footprints and higher power consumption. Because of this, some platforms [28, 21, 63] opt to use a lower power compute unit like a Raspberry Pi on the vehicles themselves for command execution and sensor data acquisition only and offload compute-intensive tasks such as control strategies and computer vision algorithms to a more powerful central computer that is connected through a network to the compute units on the vehicles. This increases the complexity of the setup as network communication has to be handled but allows for easier reconfiguration and upgrades as only the central compute unit has to be changed. This also allows to use external high performance compute clusters to handle complex control strategies.
Especially control strategies involving a camera as a sensor or strategies fusing multiple sensors require significant performance from the computing unit. Platforms that employ these strategies usually use one of the higher end options listed, namely an Nvidia Jetson or regular PCs, either on the vehicles or as a central computer unit. The smaller and the cheaper the platforms tend to be, the smaller also the compute units have to be, as size and energy constraints increase with smaller sizes. Smaller platforms also usually employ smaller and more low level sensors which can only be interfaced with through low level compute units like microcontrollers or SBCs.
III-H Vehicle-to-Everything (V2X)
V2V refers to the ability of autonomous vehicles to communicate with one another to collaboratively perform cooperative driving tasks. Its extension, Vehicle-to-Infrastructure (V2I), involves communication between vehicles and infrastructure elements, such as traffic lights or traffic signs. In 2019, the definition was expanded to V2X, which encompasses V2V and V2I, but also extends to communication with broader systems like the power grid or cellular networks. In full-scale autonomous vehicles, dedicated network components and protocols facilitate this communication.
Small-scale vehicles play a crucial role in researching these types of interactions, as they enable easy deployment of complex setups involving both vehicles and infrastructure in compact testbeds. For a small-scale platform to support such research, it must have networking capabilities, which excludes many of the educational platforms discussed in Section II-A1. Most other platforms rely on WiFi for networking, with the notable exception of the Kilobot [22], which uses a proprietary infrared communication link. WiFi enables both centralized approaches, where all vehicles and infrastructure elements communicate with a central coordinator, and decentralized approaches, where vehicles communicate directly with each other. All of the platforms investigated in our research rely on a central coordinator to manage communication, as seen in [21] and [63]. However, the open-source platforms referenced in Section II-A3 can be extended to support alternative communication strategies. Most of the platforms we examined only support V2V communication, with the Duckietown [107] and AutoDRIVE [59] platforms being exceptions. These two platforms also support V2I communication in the form of connected traffic lights. In Duckietown, the traffic lights are essentially built with the same hardware as the vehicles, allowing them to sense other vehicles through onboard cameras and accept control commands from either a central coordinator or other vehicles.
@@ IV Benchmarking
As discussed in Section III, small-scale cars primarily apply mature AV techniques from full-scale cars under simplified conditions. Thus, autonomous driving systems for small-scale vehicles are typically categorized into two distinct pipelines, similar to those used in full-scale cars: the end-to-end system pipeline and the modular system pipeline. As illustrated in Fig. 2, a modular system comprises multiple subsystems that perform various tasks such as perception and localization, mapping, path planning, and control [108, 109]. Each subsystem focuses on specific functionalities and tasks. First is the perception system, where sensors such as cameras and LiDAR are used to gather information about the surroundings, including lane markings, obstacles, traffic signs, and other vehicles. Following this, the localization and mapping system leverages GPS, odometry, IMU, or techniques like simultaneous localization and mapping (SLAM) to precisely pinpoint the position within the environment while concurrently creating a detailed map of its surroundings. The perception system creates an intermediate representation of the environment for subsequent utilization.
@@
The representation module then uses the information from the perception module and further processes the sensor data with sensor fusion techniques or creates an object map with the predicted state of each object within the sensor range. The combination of the perception module and the representation module can be seen as the scene understanding, which provides an abstract high-level representation of the environment [109]. Afterward, the planning system maps out a safe and efficient route to reach the destination. Normally, in the autonomous driving system, the planning phase is divided into two different parts, namely global path planning and local path planning [110]. Global path planning refers to the process of determining an optimal or feasible route from the current position to its destination and is done considering the entire environment and involves high-level decision-making. Local path planning or path following, on the other hand, focuses on the immediate surroundings of the vehicle and deals with making real-time adjustments to adhere to the planned global path. Finally, the control module generates driving commands based on the processed information. In control theory, the primary goal is to minimize a cost function. Various methodologies are employed for the control system, classical controllers such as the Proportional–Integral–Derivative (PID) controller, Model Predictive Control (MPC), and ML-based controllers such as IL or DRL. PID works by continuously calculating the error between a desired setpoint and a measured process variable, then applying corrective actions based on three components: proportional, integral, and derivative. The proportional component reacts to the current error, providing immediate corrections, the integral component addresses accumulated past errors to eliminate steady-state error, and the derivative component predicts future errors by considering the rate of change, enhancing stability and responsiveness. MPC, on the other hand, is an advanced control strategy designed for complex systems. It predicts future system behavior using a dynamic model and computes optimal control inputs by solving a constrained optimization problem at every time step.
For ML-based controllers, as illustrated in Fig. 3, IL involves training an agent to mimic expert behavior by learning from a dataset of demonstrations. By directly leveraging expert trajectories, IL bypasses the need for explicitly defined reward functions, making it particularly effective in scenarios where optimal behavior is well understood but challenging to formalize mathematically. Key IL algorithms include Behavioral Cloning (BC) [111], Generative Adversarial Imitation Learning (GAIL) [112], Inverse Reinforcement Learning (IRL) [113], and Dataset Aggregation (DAgger) [114]. On the other hand, RL involves training an agent to autonomously discover optimal policies by interacting with an environment and receiving feedback in the form of rewards [115]. This exploration-based approach allows RL agents to learn from scratch and adapt to a wide range of environments, even in the absence of expert guidance.
@@
An end-to-end system is characterized by a unified architecture that aims to learn the entire mapping directly from raw sensor inputs to driving actions without explicitly decomposing the task into separate modules [116, 117, 118]. It involves training a comprehensive learning model with ML methods, IL or DRL, directly processing raw sensor data and output control commands. End-to-end systems potentially simplify the system architecture by eliminating the need for handcrafted modules and feature engineering.
In this section, we focus on the two primary systems and the key tasks associated with each, as depicted in Fig. 3. Rather than delving into the detailed methodologies used to accomplish these tasks, we provide references to the relevant studies for further exploration. Additionally, we propose a baseline framework for each system, aiming to provide inspiration and guidance for researchers in developing their own approaches.
Figure 3: Classification of research in small-scale cars into modular and end-to-end systems. Modular systems are further subdivided into perception tasks, planning tasks, and control tasks.
IV-A Perception
In a modular system, the perception system performs several key functions essential to autonomous driving, including mapping the environment and localizing the autonomous vehicle, detecting lanes, and detecting objects. In this subsection, we will discuss various aspects that the perception module needs to consider.
IV-A1 Localization and Mapping